OpenStudy (anonymous):

How long does it take a train traveling at a constant rate of 70 miles per hour to travel 350 miles? 280 hours 20 hours 5 hours 245 hours

OpenStudy (alexandervonhumboldt2):

to find it, you need to divide the number of miles by number of miles per hour: 350/70=5 thus 5 hours. remember: \[s=v*t\] where `s` is distance, `v` is velocity and `t` is time.

OpenStudy (anonymous):

umm walk me thru it plz

OpenStudy (alexandervonhumboldt2):

ok: the basic formula is s=v*t where s is distance, v is velocity and t is time. in your case you don't know t, but you know v=70 mph and s=350. substitute into formula: 350=70*t t=350/70=5 hours

OpenStudy (anonymous):

5 hours ok thats the answer?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

aww thank you i am trying to learn stuff after i forgot every thing in a car accident

OpenStudy (anonymous):

im in 8 grade